# and quantum information science

### v2

@@ 6.1.1 Recursion method and tridiagonalization
where $\mathcal{K}$ is the dimension of the span of the time-evolving state, $\lvert K_{0}\rangle=\lvert\psi_{0}\rangle$, and the remaining $\lvert K_{n}\rangle$ are constructed by the iterative Lanczos algorithm [141]:
@@
In other other words, the Hamiltonian in this basis can be written as a matrix with entries $a_{n}$ along the diagonal, and entries $b_{n}$ just above and below it. For finite dimensional systems, this is called the Hessenberg form of the Hamiltonian. Numerically stable algorithms for computing the Hessenberg form of a matrix are implemented by standard libraries in Python and Mathematica (see [22]). These methods improve upon the numerically unstable Lanczos procedure, but generally require a change of basis to represent the desired initial state in a canonical form required by the implemented algorithm. We can use the Hessenberg form to extract the Lanczos coefficients without explicitly constructing the Krylov basis.
@@ 6.1.2 Moment method and the survival amplitude
To see why, suppose that a quantum system is localized, at $t=0$, in the initial state $\lvert K_{0}\rangle$. The Markov-like structure of Eq. (62) means that the probability amplitude for remaining in $\lvert K_{0}\rangle$ after one application of the Hamilton is $-ia_{0}$, while the amplitude for transitioning to $\lvert K_{1}\rangle$ is $-ib_{1}$. Next, let us apply the Hamiltonian a second time. The amplitude to wind up in $\lvert K_{0}\rangle$ again is $i^{2}(a_{0}^{2}+b_{1}^{2})$. But we already know $a_{0}$ from the first step. So, if we we also know the second moment of the Hamiltonian $\langle K_{0}\rvert H^{2}\lvert K_{0}\rangle$ through some means, then we can extract $b_{0}$ by subtracting $a_{0}^{2}$ and taking a square root. Proceeding recursively in this way, we can show that each successive moment allows the extraction of an additional Lanczos coefficient – the $a_{n}$ are computed from the odd moments by dividing out products of already computed $b_{n}$s, and the $b_{n}$s are likewise computed from the even moments. Applied numerically, this procedure is potentially sensitive to rounding errors from repeated divisions, but there are methods for circumventing this potential obstacle [22]. [^19] One possible way to resolve this difficulty, e.g., is to find squares of the Lanczos coefficients by using rational arithmetics [142]. We can use the moment method to compute the Lanczos coefficients without constructing the Krylov basis.
@@ 6.1.3 Survival amplitude
$S(t)\equiv\langle\psi(t)|\psi_{0}\rangle=\langle\psi_{0}|e^{-iHt}|\psi_{0}\rangle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mu_{n}=\langle K_{0}\rvert(-iH)^{n}\lvert K_{0}\rangle=\frac{d^{n}}{dt^{n}}S(t)\Big{|}_{t=0}\,.$ (65)
@@
which has been used to study late time quantum chaos [143, 144] in random matrix theory and quantum gravity, quantum speed limits [145], and other phenomena that depend on energy spectrum statistics.
@@ 6.1.4 Coarse-grained Lanczos spectrum and the density of states
Suppose we tridiagonalize the Hamiltonian by the recursion method or the moment method, as described above. The resulting Hamiltonian, written entirely in terms of the Lanczos coefficients, should still have the same spectrum as the original one. So we may wonder whether we can directly relate the energy spectrum, or equivalently the density of states, to the tridiagonal Lanczos spectrum (i.e., the distribution of Lanczos coefficients). Indeed, as we will discuss below, we can find such a relation in the limit of large system size, given some smoothness assumptions and after coarse-graining the Lanczos spectrum [23].
Consider a quantum system with an $N$ dimensional Hilbert space. Let $n$ be the Lanczos index and define $x=n/N$. In the large $N$ limit, $x$ is a real number between 00 and $1$ and it is convenient to write the Lanczos coefficients as functions $a(x)$ and $b(x)$ instead of as $a_{xN}$ and $b_{xN}$. Let us consider Hamiltonians and initial states for which $a(x)$ and $b(x)$ have a continuous large $N$ limit. As we will explain, if the large $N$ limit is continuous in this way, there is an analytical formula relating the density states of the theory to the Lanczos spectrum, coarse-grained in a way that we describe below.
@@
[…]
This analytical formula universally relates the density of states and the coarse-grained Lanczos coefficients for large systems. For the systems like the SYK model and Random Matrix Theories that are defined by ensembles of Hamiltonians, this formula will also relate the average density of states to the ensemble average of Lanczos coefficients for random initial states. The authors of [23] give a more formal analysis for why this formula is true, and explain how to invert it to calculate $a(x)$ and $b(x)$ from the density of states. The formula is valid for initial states for which the large $N$ limit of $a(x)$ and $b(x)$ is continuous, and so long as $x$ is not too close to the edge, i.e., $x\sim O(1/N)$. At the edge, we have to use the recursion or moment methods described above to find the Lanczos coefficients step by step.
@@ 6.1.5 Defining spread complexity
From the 1d chain perspective, the wavefunction starts out localized on $\lvert K_{0}\rangle$. As time passes, the peak of the wavefunction moves to the right (higher Krylov index $n$) and spreads out. The wavefunction has a sharp, tsunami-like edge as it moves outward [22]. This recalls the sharp edge seen in the spread of entanglement after a quench [146, 147, 148, 149, 150], at least in two-dimensional conformal systems and in theories with a holographic dual. In chaotic systems like Random Matrix Models (RMTs), the spreading wavefunction retains this coherence until it “bounces off” the end of the Krylov chain ($n=\mathcal{K}$) and settles down to equilibrium with broad support on all the basis elements. This bounce arises distinctively from the spectral correlations in chaotic theories. This was illustrated in [22] by comparing chaotic RMTs in the Gaussian Unitary, Orthogonal, and Symplectic Ensembles (GUE, GOE, GSE) with systems having the same density of states but Poisson, i.e., uncorrelated, spectral statistics.
Complete information about the spread of the wavefunction is contained in the distribution (72). We could quantify the width of this distribution in various ways. For example, we could consider moments of the distribution $\langle n^{s}\rangle=\sum_{n}n^{s}\,p_{n}(t)$. The first moment, or average position in the Krylov chain, was called the spread complexity in [22]
@@
and the analogous quantity for the spread of operators (which we will discuss below) is often referred to as the Krylov complexity (its study was initiated in [21]). These quantities have proved useful in characterizing many aspects of quantum dynamics and, as we will discuss in Sec. 7, turn out to be geometrized in the duality between the SYK model and Jackiw-Teitelboim gravity as the wormhole’s length.
@@
While the spread complexity in (73) has proven useful in many applications, alternative measures of wavefunction spread surely have their uses as well. First of all there are the higher moments we mentioned above – see, e.g., [151]. Another natural measure of spread would be the entropy of the distribution (72) or its exponential
$H(t)=-\sum_{n=1}^{\mathcal{K}-1}p_{n}(t)\ln p_{n}(t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{C}_{H}(t)\equiv e^{H(t)}\,,$ (74)
@@
This quantity was studied in the operator and state contexts in [152, 153, 154, 155], but has not yet been widely applied.
@@
Of course, we could have defined a distribution like (72) in any basis. The proofs in [22] demonstrated that the Krylov basis minimizes both spread complexity $\mathcal{C}_{K}(t)$ and the entropic measure $\mathcal{C}_{H}(t)$ at least in some finite time interval starting at $t=0$. The Krylov basis is also, in a sense, the most classical representation of the spreading wavefunction. This is because, as shown in [156], the Wigner function associated to the one dimensional Krylov chain has the lowest possible negativity as compared to any other basis. The Wigner function is the closest thing in quantum mechanics to a classical description of the state. Indeed, a quantum system with a Wigner function that is everywhere positive for all times can be efficiently simulated on a classical computer through standard methods for describing statistical systems. But if the Wigner distribution has negative regions, quantum phenomena play an essential role, indicating a likely advantage for quantum computers in simulating these systems.
@@
Below we will apply these methods to many physical systems including particles moving on group manifolds, spin systems, the SYK model, random matrix theory (RMT), and Jackiw-Teitelboim gravity. For lack of space we will not give a detailed discussion of applications to quantum billards [157, 158, 159], which are especially fascinating because such systems can have vanishing Lyapunov exponents, but can nevertheless manifest chaos. For example, the triangular billiards studied in [159] can be integrable, pseudointegrable or non-integrable, depending on the internal angles which control how bundles of classical trajectories do or do not diverge when bouncing off the walls and corners. As such, billiard systems provide a tractable venue for studying how integrability and non-integrability affect the spread of states in a quantum system. Billiards also give a way of studying the effects of symmetry on the spread of wavefunctions. For example, pseudointegrable and non-integrable isosceles triangular billiards have independent sectors that are symmetric and antisymmetric under reflection across the diagonal, and these sectors separately reproduce characteristics of chaotic theories, although the complete dynamics approximates some aspects of an integrable system, e.g., a Poisson distributed spectrum [159]. It is interesting to more broadly consider how symmetries and their breaking may affect the spread of the wavefunction. We will give some examples of these effects below for particles on group manifolds and different random matrix symmetry classes. The effects of discrete symmetries have been studied in [160] whose authors developed a framework for understanding the effects of parity and time-reversal symmetry, and their breaking, on the spread of wavefunctions. Among other things they found interesting localization phase transitions on the Krylov chain. In some of these systems such as the SYK model and RMTs, there are efficient techniques to calculate the average of the survival amplitude $\overline{S(t)}$ over an ensemble of theories. Note that in these cases $\overline{S(t)}\,\,\overline{S(t^{\prime})}$ need not equal $\overline{S(t)S(t^{\prime})}$. This similarly implies that correlations between Lanczos coefficients that affect the structure of the dynamics must be directly computed and will not be accessed by just computing the average of these coefficients over the pertinent ensemble.
@@
Above we focused entirely on systems with time-independent Hamiltonians. It would be interesting to extend the methods to systems with time-varying Hamiltonians. We also only discussed closed systems and pure quantum states, but many of the techniques can be extended to open systems and mixed states [161, 162], as reviewed in [3]. Since the density matrix describing mixed states can be considered an operator on the Hilbert space, it is natural to think about this in terms of the spread of operators, to which we turn now.
@@ 6.2 Krylov complexity of evolving operators
In Sec. 6.1, we followed [22] to quantify how quantum states spread dynamically over the Hilbert space. In this section, we will discuss a similar approach to the spread of operators [21], which was introduced earlier than Ref. [22]. We started with a description of the spread of states because it is technically simpler to some extent, and the proofs of minimality of the Krylov basis were developed in that context. But the relation between the operator-based methods of this subsection and the state-based methods of the previous section is analogous to the relation between the Heisenberg and Schrödinger approaches to quantum mechanics – either approach may be more convenient depending on the question we are studying.
The idea of defining operator size through an expansion in a natural basis was, to our knowledge, first explored in holographic systems such as the SYK model [163, 164, 165, 166]. This concept was later quantified in [21], whose authors studied operator spread in the Krylov basis, and proposed to quantify operator size in terms of the Krylov complexity (the operator analog of (73)). In this framework, we consider the Heisenberg evolution of an initial, “simple” operator $\mathcal{O}(0)$
@@
where in the second step we introduced the Liouvillian super-operator $\mathcal{L}=[H,\,\cdot\,]$. As before, we can expand the operator $\mathcal{O}(t)$ as a power series in $t$ that involves nested commutators of the Hamiltonian with $\mathcal{O}(0)$, or equivalently, various powers of the Liouvillian acting on $\mathcal{O}(0)$. As in Sec. 6.1, this series builds the Krylov subspace to which we want to apply the Lanczos procedure. If the Hilbert space is $N$–dimensional, the associated operator space in which we carry out this procedure is $N^{2}$–dimensional because operators have matrix elements $O_{ab}$, with $a,b=1...,N$.
The procedure for extracting the Krylov basis for operators and the associated Lanczos coefficients is the same as described in Sec. 6.1 for states, after vectorizing operators by mapping them into an auxiliary Hilbert space by the GNS (Gelfand-Naimark-Segal) construction (see, e.g., the discussion in [167]). In quantum information theory, such a vectorization also appears as the Choi-Jamiolkowski isomorphism, otherwise known as channel-state duality [168, 169]. After vectorizing operators in this way, $\mathcal{O}\to|\mathcal{O})$, we also have to choose an inner product on the auxiliary Hilbert space. The authors of [167] showed that the Krylov complexity defined with respect to the alternative choices behaves differently as a function of time. However, the approach described in Sec. 6.1 suggests that we should pick the inner product that produces a Krylov basis that minimizes the spread of operators. As discussed in [22], the results of [167] imply that this minimization is achieved by the Wightman inner product. Indeed, this was the choice advocated in [21], albeit for other reasons. Thus, we set
@@
Solving it yields the probability $p_{n}(t)=|\varphi_{n}(t)|^{2}$ and the Krylov complexity, which quantifies the operator size, is computed by (73). Interestingly, [170] showed that the dimension $\mathcal{K}$ of the Krylov space for operators is bounded by
@@
The spread of states and operators can be used as a probe for understanding the dynamics of quantum systems. Below, we will show how to use this tool in a number of different systems. General results can also be obtained. For example, the authors of the pioneering paper [21] performed numerical and analytical studies of Krylov complexity in integrable and chaotic systems, including the SYK model, and found that the growth of Lanczos coefficients $b_{n}$ is at most linear in $n$
@@
They conjectured that theories in which operator growth saturates this bound are chaotic, in the same sense as probed by the OTOC correlators (see Sec. 2 for a definition of these correlators). In particular, they related the coefficient $\alpha$ to the characteristic exponent $\lambda_{K}$ in an exponential growth of the Krylov complexity for chaotic theories, i.e., $\lambda_{K}=2\alpha$. In the case of the SYK model, this exponent coincided with the Lyapunov exponent read from the OTOC, which we label with $\lambda_{\text{OTOC}}$. In other words, they observed that for the SYK model $\lambda_{K}=\lambda_{\text{OTOC}}$. This match will be further discussed in section 6.3.3. However, in many theories, including 2D CFTs, operator return amplitudes are universal and do not readily distinguish between integrable and chaotic models. Therefore, the hypothesis of [21] requires refinement; indeed, as we will discuss below in Sec. 6.3.4, the difference between chaos and integrability may be more apparent in the cross-correlations between the Lanczos coefficients and the effects of these correlations on the dynamics.

### v3

@@ 6.1.1 Recursion method and tridiagonalization
where $\mathcal{K}$ is the dimension of the span of the time-evolving state, $\lvert K_{0}\rangle=\lvert\psi_{0}\rangle$, and the remaining $\lvert K_{n}\rangle$ are constructed by the iterative Lanczos algorithm Lanczos:1950zz:
@@
In other other words, the Hamiltonian in this basis can be written as a matrix with entries $a_{n}$ along the diagonal, and entries $b_{n}$ just above and below it. For finite dimensional systems, this is called the Hessenberg form of the Hamiltonian. Numerically stable algorithms for computing the Hessenberg form of a matrix are implemented by standard libraries in Python and Mathematica (see Balasubramanian:2022tpr). These methods improve upon the numerically unstable Lanczos procedure, but generally require a change of basis to represent the desired initial state in a canonical form required by the implemented algorithm. We can use the Hessenberg form to extract the Lanczos coefficients without explicitly constructing the Krylov basis.
@@ 6.1.2 Moment method and the survival amplitude
To see why, suppose that a quantum system is localized, at $t=0$, in the initial state $\lvert K_{0}\rangle$. The Markov-like structure of Eq. (62) means that the probability amplitude for remaining in $\lvert K_{0}\rangle$ after one application of the Hamilton is $ia_{0}$, while the amplitude for transitioning to $\lvert K_{1}\rangle$ is $ib_{1}$. Next, let us apply the Hamiltonian a second time. The amplitude to wind up in $\lvert K_{0}\rangle$ again is $i^{2}(a_{0}^{2}+b_{1}^{2})$. But we already know $a_{0}$ from the first step. So, if we we also know the second moment of the Hamiltonian $\langle K_{0}\rvert H^{2}\lvert K_{0}\rangle$ through some means, then we can extract $b_{0}$ by subtracting $a_{0}^{2}$ and taking a square root. Proceeding recursively in this way, we can show that each successive moment allows the extraction of an additional Lanczos coefficient – the $a_{n}$ are computed from the odd moments by dividing out products of already computed $b_{n}$s, and the $b_{n}$s are likewise computed from the even moments. Applied numerically, this procedure is potentially sensitive to rounding errors from repeated divisions, but there are methods for circumventing this potential obstacle Balasubramanian:2022tpr. [^22] One possible way to resolve this difficulty, e.g., is to find squares of the Lanczos coefficients by using rational arithmetics Uskov:2024xac. We can use the moment method to compute the Lanczos coefficients without constructing the Krylov basis.
@@ 6.1.3 Survival amplitude
$S(t)\equiv\langle\psi(t)|\psi_{0}\rangle=\langle\psi_{0}|e^{-iHt}|\psi_{0}\rangle\penalty 10000\ \penalty 10000\ \penalty 10000\ \Longrightarrow\penalty 10000\ \penalty 10000\ \penalty 10000\ \mu_{n}=\langle K_{0}\rvert(-iH)^{n}\lvert K_{0}\rangle=\frac{d^{n}}{dt^{n}}S(t)\Big|_{t=0}\,.$ (65)
@@
which has been used to study late time quantum chaos Guhr:1997ve, Cotler:2016fpe in random matrix theory and quantum gravity, quantum speed limits delCampo:2017bzr, and other phenomena that depend on energy spectrum statistics.
@@ 6.1.4 Coarse-grained Lanczos spectrum and the density of states
Suppose we tridiagonalize the Hamiltonian by the recursion method or the moment method, as described above. The resulting Hamiltonian, written entirely in terms of the Lanczos coefficients, should still have the same spectrum as the original one. So we may wonder whether we can directly relate the energy spectrum, or equivalently the density of states, to the tridiagonal Lanczos spectrum (i.e., the distribution of Lanczos coefficients). Indeed, as we will discuss below, we can find such a relation in the limit of large system size, given some smoothness assumptions and after coarse-graining the Lanczos spectrum Balasubramanian:2022dnj.
Consider a quantum system with an $N$ dimensional Hilbert space. Let $n$ be the Lanczos index and define $x=n/N$. In the large $N$ limit, $x$ is a real number between 0 and 11 and it is convenient to write the Lanczos coefficients as functions $a(x)$ and $b(x)$ instead of as $a_{xN}$ and $b_{xN}$. Let us consider Hamiltonians and initial states for which $a(x)$ and $b(x)$ have a continuous large $N$ limit. As we will explain, if the large $N$ limit is continuous in this way, there is an analytical formula relating the density states of the theory to the Lanczos spectrum, coarse-grained in a way that we describe below.
@@
$\rho(E)=\int_{0}^{1}dx\,\frac{\Theta(4b(x)^{2}-(E-a(x))^{2})}{\pi\sqrt{4b(x)^{2}-(E-a(x))^{2}}}\,.$ (71)
This analytical formula universally relates the density of states and the coarse-grained Lanczos coefficients for large systems. For the systems like the SYK model and Random Matrix Theories that are defined by ensembles of Hamiltonians, this formula will also relate the average density of states to the ensemble average of Lanczos coefficients for random initial states. The authors of Balasubramanian:2022dnj give a more formal analysis for why this formula is true, and explain how to invert it to calculate $a(x)$ and $b(x)$ from the density of states. The formula is valid for initial states for which the large $N$ limit of $a(x)$ and $b(x)$ is continuous, and so long as $x$ is not too close to the edge, i.e., $x\sim O(1/N)$. At the edge, we have to use the recursion or moment methods described above to find the Lanczos coefficients step by step.
@@ 6.1.5 Defining spread complexity
From the 1d chain perspective, the wavefunction starts out localized on $\lvert K_{0}\rangle$. As time passes, the peak of the wavefunction moves to the right (higher Krylov index $n$) and spreads out. The wavefunction has a sharp, tsunami-like edge as it moves outward Balasubramanian:2022tpr. This recalls the sharp edge seen in the spread of entanglement after a quench Calabrese:2009qy, Abajo-Arrastia:2010ajo, Balasubramanian:2011ur, Balasubramanian:2010ce, Liu:2013qca, at least in two-dimensional conformal systems and in theories with a holographic dual. In chaotic systems like Random Matrix Models (RMTs), the spreading wavefunction retains this coherence until it “bounces off” the end of the Krylov chain ($n=\mathcal{K}$) and settles down to equilibrium with broad support on all the basis elements. This bounce arises distinctively from the spectral correlations in chaotic theories. This was illustrated in Balasubramanian:2022tpr by comparing chaotic RMTs in the Gaussian Unitary, Orthogonal, and Symplectic Ensembles (GUE, GOE, GSE) with systems having the same density of states but Poisson, i.e., uncorrelated, spectral statistics.
Complete information about the spread of the wavefunction is contained in the distribution (72). We could quantify the width of this distribution in various ways. For example, we could consider moments of the distribution $\langle n^{s}\rangle=\sum_{n}n^{s}\,p_{n}(t)$. The first moment, or average position in the Krylov chain, was called the spread complexity in Balasubramanian:2022tpr
@@
and the analogous quantity for the spread of operators (which we will discuss below) is often referred to as the Krylov complexity (its study was initiated in Parker:2018yvk). These quantities have proved useful in characterizing many aspects of quantum dynamics and, as we will discuss in Sec. 7, turn out to be geometrized in the duality between the SYK model and Jackiw-Teitelboim gravity as the wormhole’s length.
@@
While the spread complexity in (73) has proven useful in many applications, alternative measures of wavefunction spread surely have their uses as well. First of all there are the higher moments we mentioned above – see, e.g., Caputa:2021ori. Another natural measure of spread would be the entropy of the distribution (72) or its exponential
$H(t)=-\sum_{n=1}^{\mathcal{K}-1}p_{n}(t)\ln p_{n}(t)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{C}_{H}(t)\equiv e^{H(t)}\,$, (74)
@@
This quantity was studied in the operator and state contexts in Barbon:2019wsy, Caputa:2021sib, Patramanis:2021lkx, Caputa:2025dep, but has not yet been widely applied.
@@
Of course, we could have defined a distribution like (72) in any basis. The proofs in Balasubramanian:2022tpr demonstrated that the Krylov basis minimizes both spread complexity $\mathcal{C}_{K}(t)$ and the entropic measure $\mathcal{C}_{H}(t)$ at least in some finite time interval starting at $t=0$. The Krylov basis is also, in a sense, the most classical representation of the spreading wavefunction. This is because, as shown in Basu:2024tgg, the Wigner function associated to the one dimensional Krylov chain has the lowest possible negativity as compared to any other basis. The Wigner function is the closest thing in quantum mechanics to a classical description of the state. Indeed, a quantum system with a Wigner function that is everywhere positive for all times can be efficiently simulated on a classical computer through standard methods for describing statistical systems. But if the Wigner distribution has negative regions, quantum phenomena play an essential role, indicating a likely advantage for quantum computers in simulating these systems Gottesman:1998hu, Veitch:2012ttw, Aaronson:2004xuh, Mari:2012ypq.
@@
Below we will apply these methods to many physical systems including particles moving on group manifolds, spin systems, the SYK model, random matrix theory (RMT), and Jackiw-Teitelboim gravity. For lack of space we will not give a detailed discussion of applications to quantum billards Hashimoto:2023swv, Camargo:2023eev, Balasubramanian:2024ghv, which are especially fascinating because such systems can have vanishing Lyapunov exponents, but can nevertheless manifest chaos. For example, the triangular billiards studied in Balasubramanian:2024ghv can be integrable, pseudointegrable or non-integrable, depending on the internal angles which control how bundles of classical trajectories do or do not diverge when bouncing off the walls and corners. As such, billiard systems provide a tractable venue for studying how integrability and non-integrability affect the spread of states in a quantum system. Billiards also give a way of studying the effects of symmetry on the spread of wavefunctions. For example, pseudointegrable and non-integrable isosceles triangular billiards have independent sectors that are symmetric and antisymmetric under reflection across the diagonal, and these sectors separately reproduce characteristics of chaotic theories, although the complete dynamics approximates some aspects of an integrable system, e.g., a Poisson distributed spectrum Balasubramanian:2024ghv. It is interesting to more broadly consider how symmetries and their breaking may affect the spread of the wavefunction. We will give some examples of these effects below for particles on group manifolds and different random matrix symmetry classes. The effects of discrete symmetries have been studied in Bhattacharya:2023yec whose authors developed a framework for understanding the effects of parity and time-reversal symmetry, and their breaking, on the spread of wavefunctions. Among other things they found interesting localization phase transitions on the Krylov chain. In some of these systems such as the SYK model and RMTs, there are efficient techniques to calculate the average of the survival amplitude $\overline{S(t)}$ over an ensemble of theories. Note that in these cases $\overline{S(t)}\,\,\overline{S(t^{\prime})}$ need not equal $\overline{S(t)S(t^{\prime})}$. This similarly implies that correlations between Lanczos coefficients that affect the structure of the dynamics must be directly computed and will not be accessed by just computing the average of these coefficients over the pertinent ensemble.
@@
Above we focused entirely on systems with time-independent Hamiltonians. It would be interesting to extend the methods to systems with time-varying Hamiltonians. We also only discussed closed systems and pure quantum states, but many of the techniques can be extended to open systems and mixed states Alishahiha:2022anw, Caputa:2024vrn, as reviewed in Nandy:2024htc. Since the density matrix describing mixed states can be considered an operator on the Hilbert space, it is natural to think about this in terms of the spread of operators, to which we turn now.
@@ 6.2 Krylov complexity of evolving operators
In Sec. 6.1, we followed Balasubramanian:2022tpr to quantify how quantum states spread dynamically over the Hilbert space. In this section, we will discuss a similar approach to the spread of operators Parker:2018yvk, which was introduced earlier than Ref. Balasubramanian:2022tpr. We started with a description of the spread of states because it is technically simpler to some extent, and the proofs of minimality of the Krylov basis were developed in that context. But the relation between the operator-based methods of this subsection and the state-based methods of the previous section is analogous to the relation between the Heisenberg and Schrödinger approaches to quantum mechanics – either approach may be more convenient depending on the question we are studying.
The idea of defining operator size through an expansion in a natural basis was, to our knowledge, first explored in holographic systems such as the SYK model Roberts:2014isa, Magan:2016ehs, Roberts:2018mnp, Qi:2018bje. This concept was later quantified in Parker:2018yvk, whose authors studied operator spread in the Krylov basis, and proposed to quantify operator size in terms of the Krylov complexity (the operator analog of (73)). In this framework, we consider the Heisenberg evolution of an initial, “simple” operator $\mathcal{O}(0)$
@@
where in the second step we introduced the Liouvillian super-operator $\mathcal{L}=[H,\,\cdot\,]$. As before, we can expand the operator $\mathcal{O}(t)$ as a power series in $t$ that involves nested commutators of the Hamiltonian with $\mathcal{O}(0)$, or equivalently, various powers of the Liouvillian acting on $\mathcal{O}(0)$. As in Sec. 6.1, this series builds the Krylov subspace to which we want to apply the Lanczos procedure. If the Hilbert space is NN–dimensional, the associated operator space in which we carry out this procedure is $N^{2}$–dimensional because operators have matrix elements $O_{ab}$, with a,b=1,Na,b=1...,N.
The procedure for extracting the Krylov basis for operators and the associated Lanczos coefficients is the same as described in Sec. 6.1 for states, after vectorizing operators by mapping them into an auxiliary Hilbert space by the GNS (Gelfand-Naimark-Segal) construction (see, e.g., the discussion in Magan:2020iac). In quantum information theory, such a vectorization also appears as the Choi-Jamiolkowski isomorphism, otherwise known as channel-state duality jamiolkowski1972linear, choi1975completely. After vectorizing operators in this way, $\mathcal{O}\to|\mathcal{O})$, we also have to choose an inner product on the auxiliary Hilbert space. The authors of Magan:2020iac showed that the Krylov complexity defined with respect to the alternative choices behaves differently as a function of time. However, the approach described in Sec. 6.1 suggests that we should pick the inner product that produces a Krylov basis that minimizes the spread of operators. As discussed in Balasubramanian:2022tpr, the results of Magan:2020iac imply that this minimization is achieved by the Wightman inner product. Indeed, this was the choice advocated in Parker:2018yvk, albeit for other reasons. Thus, we set
@@
Solving it yields the probability $p_{n}(t)=|\varphi_{n}(t)|^{2}$ and the Krylov complexity, which quantifies the operator size, is computed by (73). Interestingly, Rabinovici:2020ryf showed that the dimension $\mathcal{K}$ of the Krylov space for operators is bounded by
@@
The spread of states and operators can be used as a probe for understanding the dynamics of quantum systems. Below, we will show how to use this tool in a number of different systems. General results can also be obtained. For example, the authors of the pioneering paper Parker:2018yvk performed numerical and analytical studies of Krylov complexity in integrable and chaotic systems, including the SYK model, and found that the growth of Lanczos coefficients $b_{n}$ is at most linear in $n$
@@
They conjectured that theories in which operator growth saturates this bound are chaotic, in the same sense as probed by the OTOC correlators (see Sec. 2 for a definition of these correlators). In particular, they related the coefficient $\alpha$ to the characteristic exponent $\lambda_{K}$ in an exponential growth of the Krylov complexity for chaotic theories, i.e., $\lambda_{K}=2\alpha$. In the case of the SYK model, this exponent coincided with the Lyapunov exponent read from the OTOC, which we label with $\lambda_{\text{OTOC}}$. In other words, they observed that for the SYK model $\lambda_{K}=\lambda_{\text{OTOC}}$. This match will be further discussed in section 6.3.3. However, in many theories, including 2D CFTs, operator return amplitudes are universal and do not readily distinguish between integrable and chaotic models. Therefore, the hypothesis of Parker:2018yvk requires refinement; indeed, as we will discuss below in Sec. 6.3.4, the difference between chaos and integrability may be more apparent in the cross-correlations between the Lanczos coefficients and the effects of these correlations on the dynamics.